What can be said about the domain of the function:
f \circ g
where
f(y)= \frac{4}{y-2}
and
g(x)= \frac{5}{3x-1}
Express it in terms of a union of intervals of real numbers?
obtain the graph of f , g , and f \circ g .
1
Expert's answer
2020-09-30T18:54:58-0400

Here, "f(y)=\dfrac{4}{y-2}\;and\;g(x)=\dfrac{5}{3x-1}.\\ \\"


"\mathbf{To\;find\;: Domain \;of\;f \circ g.}"


"\mathbf{Now,\;(f\circ g)(z)=f(g(z))=f\bigg(\dfrac{5}{3z-1}\bigg)=\dfrac{4}{\dfrac{5}{3z-1}-2} }"


"\mathbf{=\dfrac{4}{\dfrac{5-2(3z-1)}{3z-1}}=\dfrac{4(3z-1)}{5-6z+2}=\dfrac{12z-4}{7-6z}}"


"\mathbf{Clearly,\; f\circ g \; is\; defined\;on\;\R\; except\;when\; the\; denominator\;is\;zero,i.e.,\; when}"

"7-6z=0\;\implies z=\dfrac{7}{6}"


"\mathbf{So,\;domain\;of\;f\circ g=\R-\{{\dfrac{7}{6}}\}=(-\infty,\dfrac{7}{6})\cup(\dfrac{7}{6},-\infty)}"


Graph of f :-




Graph of g :-




Graph of f o g :-